If a farmer waits one week to sell his corn, there is a 50% chance that they will earn an extra $10 000. However, there is also a 10% chance that they will lose $30 000. Should he sell now or wait a week? Use the results of at least 25 experimental trials to support your answer.


1
Expert's answer
2021-01-14T20:12:42-0500

"P(extra~ earn = -10000) = 0.1\\
P(extra ~earn = 30000) = 0.5\\
P(extra ~earn= 0) = 1-0.1-0.5=0.4\\
E(extra ~earn) =-10000*0.1+30000*0.5+0*0.4 = 14000"

The expected extra earn of waiting one week is $14000 is more than the extra earn of zero now. Hence he should wait a week.


In MatLab

>> p = rand(1,25);

extraearn = -10000*(p<=0.1)+30000*(p>0.5);

E = mean(extraearn)


E =


    15200